# GRAND: Graph Release with Assured Node Differential Privacy

### v2

@@ B.1 Crucial tools
$\sup_{\begin{subarray}{c}u,v\in[a,b]\\ |u-v|<\delta\end{subarray}}|\hat{F}_{m}^{-1}(u)-\hat{F}_{m}^{-1}(v)|<\varepsilon.$
@@
$P\left(\sup_{\begin{subarray}{c}u,v\in[a,b]\\ |u-v|<\delta\end{subarray}}|\hat{F}_{m}^{-1}(u)-\hat{F}_{m}^{-1}(v)|\geq\varepsilon\right)\leq P(\mathscr{B}_{m}^{c})<\frac{\eta}{2}<\eta.$
@@
  • For each fixed $u\in\mathcal{U}$, the conditional CDF $F^{\,d\mid 1:(d-1)}(\cdot|u)$ is strictly increasing.
  • The conditional CDF $F^{\,d\mid 1:(d-1)}(x|u)$ is jointly continuous in $(x,u)$ for $x\in{\bf R}$ and $u\in\mathcal{U}$.
@@
Then for any closed interval $a,b]\subset(0,1)$, and any compact set $K\subset{\bf R}^{d-1}$, we have the following conclusions.
@@
    $\sup_{u_{0}\in K}\sup_{q\in[a,b]}|(\hat{F}_{m}^{\,d\mid 1:(d-1)})^{-1}(q|u_{0})-(F^{\,d\mid 1:(d-1)})^{-1}(q|u_{0})|\xrightarrow{P}0\quad\text{as }m\to\infty.$
@@
    $P\left(\sup_{q\in[a,b]}|(\hat{F}_{m}^{\,d\mid 1:(d-1)})^{-1}(q|u_{0})-(\hat{F}_{m}^{\,d\mid 1:(d-1)})^{-1}(q|u^{\prime}_{0})|\geq\varepsilon\right)<\eta.$
@@
We first decompose the proof into several steps. For notational simplicity, we write $F(\cdot|u_{0})$ for $F^{\,d\mid 1:(d-1)}(\cdot|u_{0})$ and $\hat{F}_{m}(\cdot|u_{0})$ for $\hat{F}_{m}^{\,d\mid 1:(d-1)}(\cdot|u_{0})$. Also, let $F^{-1}(\cdot|u_{0})$ denote $(F^{\,d\mid 1:(d-1)})^{-1}(\cdot|u_{0})$ and $\hat{F}_{m}^{-1}(\cdot|u_{0})$ denote $(\hat{F}_{m}^{\,d\mid 1:(d-1)})^{-1}(\cdot|u_{0})$.
@@
$P\left(\sup_{u_{0}\in K}\sup_{q\in[a,b]}|\hat{F}_{m}^{-1}(q|u_{0})-F^{-1}(q|u_{0})|>\varepsilon\right)\xrightarrow{m\to\infty}0.$
Since $F(\cdot|u_{0})$ is strictly increasing for each fixed $u_{0}$, its inverse $F^{-1}(\cdot|u_{0})$ is well-defined. By the joint continuity of $F(x|u_{0})$ in $(x,u_{0})$ and the compactness of $K\times[a,b]$ (when $F^{-1}$ maps this to a compact set), the function $F^{-1}(q|u_{0})$ is uniformly continuous on $a,b]\times K$. Therefore, for any $\varepsilon>0$, there exists $\delta_{q}>0$ such that for all $q_{1},q_{2}\in[a,b]$ and $u_{0}\in K$,
$|q_{1}-q_{2}|<\delta_{q}\implies|F^{-1}(q_{1}|u_{0})-F^{-1}(q_{2}|u_{0})|<\frac{\varepsilon}{2}.$
@@
$\mathscr{A}_{m}=\left\{\sup_{u_{0}\in K}\sup_{x\in{\bf R}}|\hat{F}_{m}(x|u_{0})-F(x|u_{0})|\leq\delta_{q}\right\}.$
From Lemma B.2, we have $P(\mathscr{A}_{m})\xrightarrow{m\to\infty}1$.
On $\mathscr{A}_{m}$, for all $u_{0}\in K$ and $x\in{\bf R}$,
$|\hat{F}_{m}(x|u_{0})-F(x|u_{0})|\leq\delta_{q}.$
@@
$|\hat{F}_{m}^{-1}(q|u_{0})-F^{-1}(q|u_{0})|\leq\max\left\{|F^{-1}(q-\delta_{q}|u_{0})-F^{-1}(q|u_{0})|,\ |F^{-1}(q+\delta_{q}|u_{0})-F^{-1}(q|u_{0})|\right\}<\frac{\varepsilon}{2}.$
@@
$|F^{-1}(q|u_{0})-F^{-1}(q|u^{\prime}_{0})|<\frac{\varepsilon}{3}.$ (13)
@@
$\sup_{u_{0}\in K}\sup_{q\in[a,b]}|\hat{F}_{m}^{-1}(q|u_{0})-F^{-1}(q|u_{0})|\xrightarrow{P}0\quad\text{as }m\to\infty.$
@@
$P\left(\sup_{u_{0}\in K}\sup_{q\in[a,b]}|\hat{F}_{m}^{-1}(q|u_{0})-F^{-1}(q|u_{0})|\geq\frac{\varepsilon}{3}\right)<\frac{\eta}{2}.$ (14)
@@
$\mathscr{B}_{m}=\left\{\sup_{u_{0}\in K}\sup_{q\in[a,b]}|\hat{F}_{m}^{-1}(q|u_{0})-F^{-1}(q|u_{0})|<\frac{\varepsilon}{3}\right\}.$
Then, $P(\mathscr{B}_{m})>1-\frac{\eta}{2}$ for all $m\geq M$.
@@
$|\hat{F}_{m}^{-1}(q|u_{0})-\hat{F}_{m}^{-1}(q|u^{\prime}_{0})|$ $\leq|\hat{F}_{m}^{-1}(q|u_{0})-F^{-1}(q|u_{0})|$
$\quad+|F^{-1}(q|u_{0})-F^{-1}(q|u^{\prime}_{0})|$
$\quad+|F^{-1}(q|u^{\prime}_{0})-\hat{F}_{m}^{-1}(q|u^{\prime}_{0})|$
@@
Here, the first and third terms are bounded by $\frac{\varepsilon}{3}$ due to (14), and the middle term is bounded by $\frac{\varepsilon}{3}$ due to the uniform continuity of $F^{-1}(\cdot|\cdot)$ in (13).
@@
$\sup_{\begin{subarray}{c}u_{0},u^{\prime}_{0}\in K\\ \|u_{0}-u^{\prime}_{0}\|<\delta_{u}\end{subarray}}\sup_{q\in[a,b]}|\hat{F}_{m}^{-1}(q|u_{0})-\hat{F}_{m}^{-1}(q|u^{\prime}_{0})|<\varepsilon.$
@@
$P\left(\sup_{\begin{subarray}{c}u_{0},u^{\prime}_{0}\in K\\ \|u_{0}-u^{\prime}_{0}\|<\delta_{u}\end{subarray}}\sup_{q\in[a,b]}|\hat{F}_{m}^{-1}(q|u_{0})-\hat{F}_{m}^{-1}(q|u^{\prime}_{0})|\geq\varepsilon\right)\leq P(\mathscr{B}_{m}^{c})<\frac{\eta}{2}<\eta.$
@@ B.2 Latent distribution consistency: one dimensional case
We first consider the one dimensional GRAND when all latent vectors are univariate ($d=1$). We assume that $\hat{F}_{m}$ which satisfies Corollary B.1 and Theorem B.1 is available, and we focus on the $n$ i.i.d latent variables $Z_{i}\sim F$, $i=1,\ldots,n$, for the released network, and $\hat{Z}_{i}$’s are the approximations of $Z_{i}$’s satisfying $\max_{i}\|{\hat{Z}_{i}-Z_{i}}\|\leq\delta_{m}$ for $\delta_{m}=o(m^{-\tfrac{2\alpha}{2\alpha+(d-1)}})$. Let $\hat{F}_{m}$ be a kernel-smoothed CDF estimator following Lemma B.2 for $d=1$.
@@
$\tilde{Z}_{i}=\hat{F}_{m}^{-1}\left(G\left(\hat{F}_{m}(\hat{Z}_{i})+e_{i}\right)\right)$
where $e_{i}$ is a Laplace random variable for the privacy budget, independent of everything else, and $G$ is the CDF of the distribution of $U(0,1)+\text{Laplace}$. We want to show that $\tilde{Z}_{i}$ follows $F$ asymptotically.
@@
$\bar{Z}_{i}=\hat{F}_{m}^{-1}\left(G\left(\hat{F}_{m}(Z_{i})+e_{i}\right)\right)\xrightarrow{d}Y\sim F\quad\text{as }m\to\infty.$
@@
$\sup_{x\in{\bf R}}|\hat{F}_{m}(x)-F(x)|\xrightarrow{P}0\quad\text{as }m\to\infty.$
@@
$\hat{F}_{m}(Z_{i})+e_{i}\xrightarrow{d}U_{i}+e_{i}\sim U(0,1)+\mathrm{Laplace}(1/\eta).$
@@
$\bar{Z}_{i}=\hat{F}_{m}^{-1}\left(G\left(\hat{F}_{m}(Z_{i})+e_{i}\right)\right)\xrightarrow{d}Y\sim F\quad\text{as }m\to\infty,$
@@
$\bar{Z}_{i}-T_{i}$ $=$ $\left[\hat{F}_{m}^{-1}\left(G\left(\hat{F}_{m}(Z_{i})+e_{i}\right)\right)-\hat{F}_{m}^{-1}\left(G\left(F(Z_{i})+e_{i}\right)\right)\right]$
$+\left[\hat{F}_{m}^{-1}\left(G\left(F(Z_{i})+e_{i}\right)\right)-F^{-1}\left(G\left(F(Z_{i})+e_{i}\right)\right)\right].$
@@
$A_{i}=\hat{F}_{m}^{-1}\left(G\left(\hat{F}_{m}(Z_{i})+e_{i}\right)\right)-\hat{F}_{m}^{-1}\left(G\left(F(Z_{i})+e_{i}\right)\right),$
$B_{i}=\hat{F}_{m}^{-1}\left(G\left(F(Z_{i})+e_{i}\right)\right)-F^{-1}\left(G\left(F(Z_{i})+e_{i}\right)\right).$
@@
$|A_{i}|=\left|\hat{F}_{m}^{-1}\left(G\left(\hat{F}_{m}(Z_{i})+e_{i}\right)\right)-\hat{F}_{m}^{-1}\left(G\left(F(Z_{i})+e_{i}\right)\right)\right|\xrightarrow{P}0.$
@@
$P\left(|\hat{F}_{m}(Z_{i})-F(Z_{i})|\geq\delta\right)\xrightarrow{m\to\infty}0.$
As $G$ is continuous on $[0,1]$, it is uniformly continuous. Therefore, for any $\varepsilon^{\prime}>0$, there exists $\delta>0$ such that for all $u,v\in[0,1]$:
@@
Thus, when $|\hat{F}_{m}(Z_{i})-F(Z_{i})|<\delta$, we have:
$|G(\hat{F}_{m}(Z_{i})+e_{i})-G(F(Z_{i})+e_{i})|<\varepsilon^{\prime}.$
@@
$P\left(\sup_{\begin{subarray}{c}u,v\in[a,b]\\ |u-v|<\delta^{\prime}\end{subarray}}|\hat{F}_{m}^{-1}(u)-\hat{F}_{m}^{-1}(v)|\geq\varepsilon\right)<\eta.$
@@
$\mathscr{E}_{m}=\left\{|\hat{F}_{m}(Z_{i})-F(Z_{i})|<\delta\right\}\cap\left\{\sup_{\begin{subarray}{c}u,v\in[a,b]\\ |u-v|<\delta^{\prime}\end{subarray}}|\hat{F}_{m}^{-1}(u)-\hat{F}_{m}^{-1}(v)|<\varepsilon\right\}.$
Since $\hat{F}_{m}(Z_{i})\xrightarrow{P}F(Z_{i})$ and the second event occurs with high probability, $P(\mathscr{E}_{m})\xrightarrow{m\to\infty}1$. Note that we can make this event even stronger as
$\mathscr{E}_{m}=\left\{\sup_{1\leq i\leq n}|\hat{F}_{m}(Z_{i})-F(Z_{i})|<\delta\right\}\cap\left\{\sup_{\begin{subarray}{c}u,v\in[a,b]\\ |u-v|<\delta^{\prime}\end{subarray}}|\hat{F}_{m}^{-1}(u)-\hat{F}_{m}^{-1}(v)|<\varepsilon\right\}$
@@
$|\hat{F}_{m}(Z_{i})-F(Z_{i})|<\delta\implies|G(u_{1})-G(u_{2})|<\varepsilon^{\prime}.$
@@
$|A_{i}|=\left|\hat{F}_{m}^{-1}(G(u_{1}))-\hat{F}_{m}^{-1}(G(u_{2}))\right|<\varepsilon.$
[…]
@@
$|B_{i}|=\left|\hat{F}_{m}^{-1}\left(G\left(F(Z_{i})+e_{i}\right)\right)-F^{-1}\left(G\left(F(Z_{i})+e_{i}\right)\right)\right|\xrightarrow{P}0.$
@@
$P\left(\sup_{u\in[a,b]}|\hat{F}_{m}^{-1}(u)-F^{-1}(u)|\geq\varepsilon\right)<\eta^{\prime}.$ (15)
@@
$P\left(U_{i}\in[a,b]\right)>1-\eta^{\prime}.$ (16)
On the event $\mathscr{E}_{m}^{\prime}=\left\{U_{i}\in[a,b]\right\}\cap\left\{\sup_{u\in[a,b]}|\hat{F}_{m}^{-1}(u)-F^{-1}(u)|<\varepsilon\right\}$, we have
@@
$P\left(|B_{i}|\geq\varepsilon\right)\leq P(\mathscr{E}_{m}^{{}^{\prime}c})\leq P\left(U_{i}\notin[a,b]\right)+P\left(\sup_{u\in[a,b]}|\hat{F}_{m}^{-1}(u)-F^{-1}(u)|\geq\varepsilon\right).$
@@
$P\left(|B_{i}|\geq\varepsilon\right)<\eta^{\prime}+\eta^{\prime}=2\eta^{\prime}.$
Since $\eta^{\prime}>0$ is arbitrary, we can make $P\left(|B_{i}|\geq\varepsilon\right)$ as small as desired by choosing $\eta^{\prime}$ appropriately and ensuring $m\geq M$. Therefore,
$|B_{i}|\xrightarrow{P}0.$
@@
$\bar{Z}_{i}-T_{i}\xrightarrow{P}0\quad\text{as }m\to\infty.$
@@
$T_{i}=F^{-1}\left(G\left(F(Z_{i})+e_{i}\right)\right)\sim F^{-1}(U^{\prime})\sim F,$
where $U^{\prime}\sim\text{Uniform}(0,1)$. Since $\bar{Z}_{i}-T_{i}\xrightarrow{P}0$ and $T_{i}\xrightarrow{d}Y\sim F$, by Slutsky’s theorem,
@@
$\tilde{Z}_{i}=\hat{F}_{m}^{-1}\left(G\left(\hat{F}_{m}(\hat{Z}_{i})+e_{i}\right)\right).$
@@
Suppose the conditions of Lemma B.3 hold. In addition, assume that the perturbation size $\delta_{m}$ satisfies $\delta_{m}\to 0$ as $m\to\infty$, and for the current set of observations $\{Z_{j}\}_{j=1}^{n}$ and their perturbed versions $\{\hat{Z}_{j}\}_{j=1}^{n}$, we have $\sup_{1\leq j\leq n}|\hat{Z}_{j}-Z_{j}|\leq\delta_{m}$ with probability tending to 1 as $m\to\infty$. Then for any given $i\in[n]$, $\tilde{Z}_{i}$ weakly converges to $F$ as $m\to\infty$.

### v3

@@ B.1 Crucial tools
$\sup_{\begin{subarray}{c}u,v\in[a,b]\\ |u-v|<\delta\end{subarray}}\big|\hat{F}_{m}^{-1}(u)-\hat{F}_{m}^{-1}(v)\big|<\varepsilon.$
@@
$\mathbb{P}\left(\sup_{\begin{subarray}{c}u,v\in[a,b]\\ |u-v|<\delta\end{subarray}}\big|\hat{F}_{m}^{-1}(u)-\hat{F}_{m}^{-1}(v)\big|\geq\varepsilon\right)\leq P(\mathscr{B}_{m}^{c})<\frac{\eta}{2}<\eta.$
@@
  • For each fixed $u\in\mathcal{U}$, the conditional CDF $F^{d\,|\,1:(d-1)}(\cdot\,|\,u)$ is strictly increasing.
  • The conditional CDF $F^{d\,|\,1:(d-1)}(x\,|\,u)$ is jointly continuous in $(x,u)$ for $x\in\mathbb{R}$ and $u\in\mathcal{U}$.
@@
Then for any closed interval $a,b]\subset(0,1)$, and any compact set $K\subset\mathbb{R}^{d-1}$, we have the following conclusions.
@@
    $\sup_{u_{0}\in K}\sup_{q\in[a,b]}\big|(\hat{F}_{m}^{d\,|\,1:(d-1)})^{-1}(q\,|\,u_{0})-(F^{d\,|\,1:(d-1)})^{-1}(q\,|\,u_{0})\big|\xrightarrow{\mathbb{P}}0\quad\text{as }m\to\infty.$
@@
    $\mathbb{P}\left(\sup_{q\in[a,b]}\big|(\hat{F}_{m}^{d\,|\,1:(d-1)})^{-1}(q\,|\,u_{0})-(\hat{F}_{m}^{d\,|\,1:(d-1)})^{-1}(q\,|\,u^{\prime}_{0})\big|\geq\varepsilon\right)<\eta.$
@@
[…]
@@
$\mathbb{P}\left(\sup_{u_{0}\in K}\sup_{q\in[a,b]}\big|\hat{F}_{m}^{-1}(q\,|\,u_{0})-F^{-1}(q\,|\,u_{0})\big|>\varepsilon\right)\xrightarrow{m\to\infty}0.$
Since $F(\cdot\,|\,u_{0})$ is strictly increasing for each fixed $u_{0}$, its inverse $F^{-1}(\cdot\,|\,u_{0})$ is well-defined. By the joint continuity of $F(x\,|\,u_{0})$ in $(x,u_{0})$ and the compactness of $a,b]\times K$ (when $F^{-1}$ maps this to a compact set), the function $F^{-1}(q\,|\,u_{0})$ is uniformly continuous on $a,b]\times K$. Therefore, for any $\varepsilon>0$, there exists $\delta_{q}>0$ such that for all $q_{1},q_{2}\in[a,b]$ and $u_{0}\in K$,
$|q_{1}-q_{2}|<\delta_{q}\implies\big|F^{-1}(q_{1}\,|\,u_{0})-F^{-1}(q_{2}\,|\,u_{0})\big|<\frac{\varepsilon}{2}.$
@@
$\mathscr{A}_{m}=\left\{\sup_{u_{0}\in K}\sup_{x\in\mathbb{R}}\big|\hat{F}_{m}(x\,|\,u_{0})-F(x\,|\,u_{0})\big|\leq\delta_{q}\right\}.$
From Lemma B.2, we have $\mathbb{P}(\mathscr{A}_{m})\xrightarrow{m\to\infty}1$.
On $\mathscr{A}_{m}$, for all $u_{0}\in K$ and $x\in\mathbb{R}$,
$\big|\hat{F}_{m}(x\,|\,u_{0})-F(x\,|\,u_{0})\big|\leq\delta_{q}.$
@@
$\big|\hat{F}_{m}^{-1}(q\,|\,u_{0})-F^{-1}(q\,|\,u_{0})\big|\leq\max\left\{\big|F^{-1}(q-\delta_{q}\,|\,u_{0})-F^{-1}(q\,|\,u_{0})\big|,\,\big|F^{-1}(q+\delta_{q}\,|\,u_{0})-F^{-1}(q\,|\,u_{0})\big|\right\}<\frac{\varepsilon}{2}.$
@@
$\big|F^{-1}(q\,|\,u_{0})-F^{-1}(q\,|\,u^{\prime}_{0})\big|<\frac{\varepsilon}{3}.$ (13)
@@
$\sup_{u_{0}\in K}\sup_{q\in[a,b]}\big|\hat{F}_{m}^{-1}(q\,|\,u_{0})-F^{-1}(q\,|\,u_{0})\big|\xrightarrow{\mathbb{P}}0\quad\text{as }m\to\infty.$
@@
$\mathbb{P}\left(\sup_{u_{0}\in K}\sup_{q\in[a,b]}\big|\hat{F}_{m}^{-1}(q\,|\,u_{0})-F^{-1}(q\,|\,u_{0})\big|\geq\frac{\varepsilon}{3}\right)<\frac{\eta}{2}.$ (14)
@@
$\mathscr{B}_{m}=\left\{\sup_{u_{0}\in K}\sup_{q\in[a,b]}\big|\hat{F}_{m}^{-1}(q\,|\,u_{0})-F^{-1}(q\,|\,u_{0})\big|<\frac{\varepsilon}{3}\right\}.$
Then, $P(\mathscr{B}_{m})>1-\eta/2$ for all $m\geq M$.
@@
$\big|\hat{F}_{m}^{-1}(q\,|\,u_{0})-\hat{F}_{m}^{-1}(q\,|\,u^{\prime}_{0})\big|$ $\leq\big|\hat{F}_{m}^{-1}(q\,|\,u_{0})-F^{-1}(q\,|\,u_{0})\big|$
$\quad+\big|F^{-1}(q\,|\,u_{0})-F^{-1}(q\,|\,u^{\prime}_{0})\big|$
$\quad+\big|F^{-1}(q\,|\,u^{\prime}_{0})-\hat{F}_{m}^{-1}(q\,|\,u^{\prime}_{0})\big|$
@@
Here, the first and third terms are bounded by $\varepsilon/3$ due to (14), and the middle term is bounded by $\varepsilon/3$ due to the uniform continuity of $F^{-1}(\cdot\,|\,\cdot)$ in (13).
@@
$\sup_{\begin{subarray}{c}u_{0},u^{\prime}_{0}\in K\\ \|u_{0}-u^{\prime}_{0}\|<\delta_{u}\end{subarray}}\sup_{q\in[a,b]}\big|\hat{F}_{m}^{-1}(q\,|\,u_{0})-\hat{F}_{m}^{-1}(q\,|\,u^{\prime}_{0})\big|<\varepsilon.$
@@
$\mathbb{P}\left(\sup_{\begin{subarray}{c}u_{0},u^{\prime}_{0}\in K\\ \|u_{0}-u^{\prime}_{0}\|<\delta_{u}\end{subarray}}\sup_{q\in[a,b]}\big|\hat{F}_{m}^{-1}(q\,|\,u_{0})-\hat{F}_{m}^{-1}(q\,|\,u^{\prime}_{0})\big|\geq\varepsilon\right)\leq\mathbb{P}(\mathscr{B}_{m}^{c})<\frac{\eta}{2}<\eta.$
@@ B.2 Latent distribution consistency: one dimensional case
We first consider the one dimensional GRAND when all latent vectors are univariate ($d=1$). We assume that $\hat{F}_{m}$ which satisfies Corollary B.1 and Theorem B.1 is available, and we focus on the $n$ i.i.d. latent variables $Z_{i}\sim F$, $i=1,\ldots,n$, for the released network, and $\hat{Z}_{i}$’s are the approximations of $Z_{i}$’s satisfying $\max_{i\in[n]}\|{\hat{Z}_{i}-Z_{i}}\|\leq\delta_{m}$ for $\delta_{m}=o\left(m^{-\tfrac{2\alpha}{2\alpha+(d-1)}}\right)$. Let $\hat{F}_{m}$ be a kernel-smoothed CDF estimator following Lemma B.2 for $d=1$.
@@
$\tilde{Z}_{i}=\hat{F}_{m}^{-1}(G(\hat{F}_{m}(\hat{Z}_{i})+e_{i}))$
where $e_{i}$ is a Laplace random variable for the privacy budget, independent of everything else, and $G$ is the CDF of the distribution of $\text{Uniform}(0,1)+\text{Laplace}(0,1/\varepsilon)$. We want to show that $\tilde{Z}_{i}$ follows $F$ asymptotically.
@@
$\bar{Z}_{i}=\hat{F}_{m}^{-1}(G(\hat{F}_{m}(Z_{i})+e_{i}))\xrightarrow{d}Y\sim F\quad\text{as }m\to\infty.$
@@
$\sup_{x\in\mathbb{R}}\big|\hat{F}_{m}(x)-F(x)\big|\xrightarrow{\mathbb{P}}0\quad\text{as }m\to\infty.$
@@
$\hat{F}_{m}(Z_{i})+e_{i}\xrightarrow{d}U_{i}+e_{i}\sim\text{Uniform}(0,1)+\text{Laplace}(0,1/\varepsilon).$
@@
$\bar{Z}_{i}=\hat{F}_{m}^{-1}(G(\hat{F}_{m}(Z_{i})+e_{i}))\xrightarrow{d}Y\sim F\quad\text{as }m\to\infty,$
@@
$\bar{Z}_{i}-T_{i}$ $=\left[\hat{F}_{m}^{-1}(G(\hat{F}_{m}(Z_{i})+e_{i}))-\hat{F}_{m}^{-1}(G(F(Z_{i})+e_{i}))\right]$
$\quad+\left[\hat{F}_{m}^{-1}(G(F(Z_{i})+e_{i}))-F^{-1}(G(F(Z_{i})+e_{i}))\right].$
@@
$A_{i}=\hat{F}_{m}^{-1}(G(\hat{F}_{m}(Z_{i})+e_{i}))-\hat{F}_{m}^{-1}(G(F(Z_{i})+e_{i})),$
$B_{i}=\hat{F}_{m}^{-1}(G(F(Z_{i})+e_{i}))-F^{-1}(G(F(Z_{i})+e_{i})).$
@@
$|A_{i}|=\big|\hat{F}_{m}^{-1}(G(\hat{F}_{m}(Z_{i})+e_{i}))-\hat{F}_{m}^{-1}(G(F(Z_{i})+e_{i}))\big|\xrightarrow{\mathbb{P}}0.$
@@
$\mathbb{P}\left(\big|\hat{F}_{m}(Z_{i})-F(Z_{i})\big|\geq\delta\right)\xrightarrow{m\to\infty}0.$
As $G$ is uniformly continuous, for any $\varepsilon^{\prime}>0$, there exists $\delta>0$ such that
@@
Thus, when $\big|\hat{F}_{m}(Z_{i})-F(Z_{i})\big|<\delta$, we have:
$\big|G(\hat{F}_{m}(Z_{i})+e_{i})-G(F(Z_{i})+e_{i})\big|<\varepsilon^{\prime}.$
@@
$\mathbb{P}\left(\sup_{\begin{subarray}{c}u,v\in[a,b]\\ |u-v|<\delta^{\prime}\end{subarray}}\big|\hat{F}_{m}^{-1}(u)-\hat{F}_{m}^{-1}(v)\big|\geq\varepsilon\right)<\eta.$
@@
$\mathscr{E}_{m}=\left\{\big|\hat{F}_{m}(Z_{i})-F(Z_{i})\big|<\delta\right\}\cap\left\{\sup_{\begin{subarray}{c}u,v\in[a,b]\\ |u-v|<\delta^{\prime}\end{subarray}}\big|\hat{F}_{m}^{-1}(u)-\hat{F}_{m}^{-1}(v)\big|<\varepsilon\right\}.$
Since $\hat{F}_{m}(Z_{i})\xrightarrow{\mathbb{P}}F(Z_{i})$ and the second event occurs with high probability, $\mathbb{P}(\mathscr{E}_{m})\xrightarrow{m\to\infty}1$. Note that we can make this event even stronger as
$\mathscr{E}_{m}=\left\{\sup_{1\leq i\leq n}\big|\hat{F}_{m}(Z_{i})-F(Z_{i})\big|<\delta\right\}\cap\left\{\sup_{\begin{subarray}{c}u,v\in[a,b]\\ |u-v|<\delta^{\prime}\end{subarray}}\big|\hat{F}_{m}^{-1}(u)-\hat{F}_{m}^{-1}(v)\big|<\varepsilon\right\}$
@@
$\big|\hat{F}_{m}(Z_{i})-F(Z_{i})\big|<\delta\implies|G(u_{1})-G(u_{2})|<\varepsilon^{\prime}.$
@@
$|A_{i}|=\big|\hat{F}_{m}^{-1}(G(u_{1}))-\hat{F}_{m}^{-1}(G(u_{2}))\big|<\varepsilon.$
Since $\mathbb{P}(\mathscr{E}_{m}^{c})\xrightarrow{m\to\infty}0$, we have $|A_{i}|\xrightarrow{\mathbb{P}}0.$
@@
$|B_{i}|=\big|\hat{F}_{m}^{-1}(G(F(Z_{i})+e_{i}))-F^{-1}(G(F(Z_{i})+e_{i}))\big|\xrightarrow{\mathbb{P}}0.$
@@
$\mathbb{P}\left(\sup_{u\in[a,b]}\big|\hat{F}_{m}^{-1}(u)-F^{-1}(u)\big|\geq\varepsilon\right)<\eta^{\prime}.$ (15)
@@
$\mathbb{P}(U_{i}\in[a,b])>1-\eta^{\prime}.$ (16)
On the event $\mathscr{E}_{m}^{\prime}=\left\{U_{i}\in[a,b]\right\}\cap\left\{\sup_{u\in[a,b]}\big|\hat{F}_{m}^{-1}(u)-F^{-1}(u)\big|<\varepsilon\right\}$, we have
@@
$\mathbb{P}(|B_{i}|\geq\varepsilon)\leq\mathbb{P}(\mathscr{E}_{m}^{{}^{\prime}c})\leq\mathbb{P}(U_{i}\notin[a,b])+\mathbb{P}\left(\sup_{u\in[a,b]}\big|\hat{F}_{m}^{-1}(u)-F^{-1}(u)\big|\geq\varepsilon\right).$
@@
$\mathbb{P}(|B_{i}|\geq\varepsilon)<\eta^{\prime}+\eta^{\prime}=2\eta^{\prime}.$
Since $\eta^{\prime}>0$ is arbitrary, we can make $\mathbb{P}(|B_{i}|\geq\varepsilon)$ as small as desired by choosing $\eta^{\prime}$ appropriately and ensuring $m\geq M$. Therefore,
$|B_{i}|\xrightarrow{\mathbb{P}}0.$
@@
$\bar{Z}_{i}-T_{i}\xrightarrow{\mathbb{P}}0\quad\text{as }m\to\infty.$
@@
$T_{i}=F^{-1}(G(F(Z_{i})+e_{i}))=F^{-1}(U_{i})\sim F,$
where $U_{i}\sim\text{Uniform}(0,1)$. Since $\bar{Z}_{i}-T_{i}\xrightarrow{\mathbb{P}}0$ and $T_{i}\xrightarrow{d}Y\sim F$, by Slutsky’s theorem,
@@
$\tilde{Z}_{i}=\hat{F}_{m}^{-1}(G(\hat{F}_{m}(\hat{Z}_{i})+e_{i})).$
@@
Suppose the conditions of Lemma B.3 hold. In addition, assume that the perturbation size $\delta_{m}$ satisfies $\delta_{m}\to 0$ as $m\to\infty$, and for the current set of observations $\{Z_{i}\}_{i=1}^{n}$ and their perturbed versions $\{\hat{Z}_{i}\}_{i=1}^{n}$, we have $\sup_{1\leq i\leq n}|\hat{Z}_{i}-Z_{i}|\leq\delta_{m}$ with probability tending to 1 as $m\to\infty$. Then for any given $i\in[n]$, $\tilde{Z}_{i}$ weakly converges to $F$ as $m\to\infty$.